OpenStudy (anonymous):

In solving the equation (x + 1)(x – 2) = 54, Eric stated that the solution would be x + 1 = 54 => x = 53 or (x – 2) = 54 => x = 56 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

(x+1)(x-2)=54 use FOIL to x^2 - x -2 =54 so x^2 - x -56 =0 x^2-8x+7x-56=0 x(x-8) +7(x-8)= (x-8)(x+7)=0 so x-8=0 or x+7=0 hence x=8 or x=-7

OpenStudy (anonymous):

he is using the zero-product property. However, the zero-product property only applies if the product is zero.